# JUHAN ARU, PHILÉMON BORDEREAU

### v1

@@ 2.3. Crossing $\mathrm{SLE}_{4}$
with the subscripts/superscripts $\mathcal{B}_{l}^{\prime},\mathcal{B}_{r}^{\prime}$ to differentiate when the operators correspond to Brownian excursions from $\mathcal{B}_{i}^{\prime}$ to $\mathcal{B}_{j}^{\prime}$, $i,j=l,r$.
We first finish the proof of the theorem before going over the proof of the proposition. When computing the average of Equation (24) with Lemma 6, the determinant term is that of
(25) $I-G_{D\setminus\eta}|_{\delta}(H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})_{\mathcal{B}_{l}^{\prime}}^{\mathcal{B}_{l}^{\prime}}-G_{D\setminus\eta}|_{\delta}(H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})_{\mathcal{B}_{r}^{\prime}}^{\mathcal{B}_{r}^{\prime}}-2G_{D\setminus\eta}|_{\delta}(H_{\partial D^{\prime}})_{\mathcal{B}_{l}^{\prime}}^{\mathcal{B}_{r}^{\prime}}.$
This can again be understood with another decomposition formula for Brownian loops. One can further decompose the expressions obtained after applying Lemma 1 by distinguishing on which side the Brownian path leaves and hits $\delta$:
$G_{D}-G_{D\setminus\eta}$ $=(H_{D\setminus\eta})_{\eta}^{\delta}(H_{D^{\prime}})_{\delta}^{\eta}G_{D}$
$=G_{D\setminus\eta}\left((H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})_{\mathcal{B}_{l}^{\prime}}^{\mathcal{B}_{l}^{\prime}}-(H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})_{\mathcal{B}_{r}^{\prime}}^{\mathcal{B}_{r}^{\prime}}-2(H_{\partial D^{\prime}})_{\mathcal{B}_{l}^{\prime}}^{\mathcal{B}_{r}^{\prime}}\right)G_{D}$
From here, Proposition 2 once again provides that the determinant term in the Gaussian integration is equal to
@@
(26) $\exp(\frac{1}{2}((\partial_{\nu_{l}}+\partial_{\nu_{r}})(g_{D^{\prime}}),G_{D}((\partial_{\nu_{l}}+\partial_{\nu_{r}})g_{D^{\prime}}))_{\delta})=\exp(-((\partial_{\nu_{l}}+\partial_{\nu_{r}})(g_{D^{\prime}}-g_{D}),G_{D}((\partial_{\nu_{l}}+\partial_{\nu_{r}})g_{D^{\prime}}))_{\delta}),$
given that the gradient of $g_{D}$ is single valued. Now we need to invert $G_{D}|_{\delta}$, i.e. the operator
$f\mapsto\int_{\delta}G_{D}(\cdot,x)f(x)\mathrm{d}x$
that from a function in $L^{2}(\delta)$ outputs a continuous function in $D$, harmonic in $D\setminus\delta$, with values 00 on $\partial D$ and whose normal derivative jumps by $f(x)$ across $\delta$ at $x$. If one now replaces $g_{D^{\prime}}$ with the multivalued function $g_{D^{\prime},m}$ that is equal to $g_{D}+2n\lambda$ in every branch in $D^{\prime}$, then $g_{D}-g_{D^{\prime},m}$ is single-valued and $g_{D}-g_{D^{\prime},m}$ satisfies the conditions above [^3] In fact the inverse of the Green’s function here is also a Neumann-type operator that computes the difference of the normal derivatives of the harmonic extensions of $f$ on both sides of $\delta$.: it follows that $G_{D}((\partial_{\nu_{l}}+\partial_{\nu_{r}})(g_{D}-g_{D^{\prime},m}))=(g_{D}-g_{D^{\prime},m}))|_{\delta}$. We rewrite Equation (26) as
$\exp(\frac{1}{2}((g_{D}-g_{D^{\prime},m}),(\partial_{\nu_{l}}+\partial_{\nu_{r}})g_{D^{\prime}})_{\delta})=\exp(\frac{1}{2}\int_{D^{\prime}}g_{D}\cdot\nabla g_{D^{\prime}}-\frac{1}{2}\norm{\nabla g_{D^{\prime}}}_{D^{\prime}}^{2})$
@@
$\int_{D^{\prime}}g_{D}\cdot\nabla g_{D^{\prime}}=\int_{\delta}g_{D^{\prime}}\cdot(\partial_{\nu_{l}}+\partial_{\nu_{r}})g_{D}=\int_{\delta}g_{D}\cdot(\partial_{\nu_{l}}+\partial_{\nu_{r}})g_{D}=\norm{\nabla g_{D}}_{D^{\prime}}^{2}$
given that $g_{D}$ has monodromy $2\lambda$. ∎
We now turn to the proof of Proposition 17.
Proof of Proposition 17.
The proof starts identically to that of Proposition 11, up to the computation of $((u_{\xi}^{\varepsilon})_{\eta},(u_{\xi}^{\varepsilon})_{\eta})_{\nabla(D^{\prime})}$. Write
@@
Similarly to the proof of Proposition 11, assume that $\xi$ is smooth, and compute
@@
$=\int_{\mathcal{B}_{l}^{\prime}}\xi\cdot(H_{\partial D^{\prime}}-H_{\partial D_{l}^{\prime}})^{\mathcal{B}_{l}^{\prime}}\xi+\int_{\mathcal{B}_{r}^{\prime}}\xi\cdot(H_{\partial D^{\prime}}-H_{\partial D_{r}^{\prime}})^{\mathcal{B}_{r}^{\prime}}\xi+2\int_{\mathcal{B}_{r}^{\prime}}\xi\cdot(H_{\partial D^{\prime}})^{\mathcal{B}_{r}^{\prime}}\xi,$
@@
The computation of the term $(\Phi_{\eta},u_{\xi}^{\varepsilon})_{\nabla(D^{\prime})}$ is identical to Equation (18), taking into account that $\mathcal{B}_{l}^{\prime}$ and $\mathcal{B}_{r}^{\prime}$ are seen as distinct parts of the boundary of $D^{\prime}$:
$(\Phi_{\eta},u_{\xi}^{\varepsilon})_{\nabla(D^{\prime})}=-\int_{\partial D^{\prime}}\xi\cdot\partial_{\nu}h_{\eta}=-\int_{\mathcal{B}_{l}^{\prime}}\xi\cdot\partial_{\nu}h_{\eta}-\int_{\mathcal{B}_{r}^{\prime}}\xi\cdot\partial_{\nu}h_{\eta}=(\partial_{\nu}g_{D^{\prime}},\xi)_{\mathcal{B}_{l}^{\prime}}+(\partial_{\nu}g_{D^{\prime}},\xi)_{\mathcal{B}_{r}^{\prime}}.$
3. Construction of $\mathrm{SLE}_{\kappa}$ in multiply connected domains
@@
We now turn to a more general construction of $\mathrm{SLE}_{\kappa}$ in multiply connected domains for the whole range $\kappa\in(8/3,4]$, arguably simpler than the previous construction of $\mathrm{SLE}_{4}$, but specific to the non-crossing case. We once again start by recalling classical facts about restriction measures and CLEs.
3.1. Restriction measures and CLEs
Consider $D\subset\mathbb{C}$ a (non-crossing) admissible domain and a Poisson point process (PPP) of intensity $h$ with respect to the (infinite) measure of Brownian excursions from $\mathcal{B}_{l}$ in $D$
$\pi\iint_{\mathcal{B}_{l}\times\mathcal{B}_{l}}\mu_{\partial D}(z,w)\mathrm{d}z\mathrm{d}w.$
Denote by $R_{h}^{D,\mathcal{B}_{l}}$ its filling, i.e. the complement of the connected component of $D$ minus a sample of the PPP containing $\mathcal{B}_{r}$. Note that the factor of $\pi$ is due to our convention for the Poisson kernel.
Lemma 18.
@@ 3.1. Restriction measures and CLEs
(27) $\exp(\frac{h\pi}{4}(\norm{u_{D^{\prime}}}_{\nabla(D^{\prime})}^{2}-\norm{u_{D}}_{\nabla(D)}^{2})),$
where $u_{D}$ (resp. $u_{D^{\prime}}$) is harmonic with boundary values $-1$ on $\mathcal{B}_{l}$, $+1$ on the rest of $\partial D$ (resp. $\partial D^{\prime}$).
@@
By definition of the PPP and the observation that the excursion measures satisfy a form of restriction (see Section 1.2), the random variable that counts the loops in a realization of the PPP that exit $D^{\prime}$ is Poisson with parameter
$h\pi\iint_{\mathcal{B}_{l}^{\prime}}(H_{\partial D}(x,y)-H_{\partial D^{\prime}}(x,y))\mathrm{d}x\mathrm{d}y.$
@@
$\frac{h\pi}{4}(\norm{u_{D^{\prime}}}_{\nabla(D^{\prime})}^{2}-\norm{u_{D}}_{\nabla(D)}^{2}).\qed$
Naturally the same construction can be replicated ad hoc replacing $\mathcal{B}_{l}$ by $\mathcal{B}_{\mathbf{b}}^{-}$ for $\mathbf{b}\in\left\{\pm\lambda\right\}^{n}$. Importantly, as mentioned in Remark 2.1 it can be verified that the regularized Dirichlet energy satisfies the following conformal covariance property, for $\varphi:D\to\varphi(D)$ a conformal equivalence:
$\exp(-\frac{1}{2}\norm{\varphi(u_{D})}_{\nabla(D),\text{reg}}^{2})=\absolutevalue{\varphi^{\prime}(x)}^{2/\pi}\absolutevalue{\varphi^{\prime}(y)}^{2/\pi}\exp(-\frac{1}{2}\norm{u_{\varphi(D)}}_{\nabla(\varphi(D)),\text{reg}}^{2}),$
so that up to a multiplicative constant,
$\exp(\norm{u_{D}}_{\nabla(D),\text{reg}}^{2})=H_{\partial D}(x,y)^{-4/\pi}$
when $D$ is simply connected. This is particularly important in the case above, as Equation (27) becomes
$\left(\frac{H_{\partial D^{\prime}}(x,y)}{H_{\partial D}(x,y)}\right)^{h},$
i.e. $R_{h}^{D,\mathcal{B}_{l}}$ is a one-sided restriction measure of exponent $h$ (see [LSW03] for more about restriction measures).

### v2

@@ 2.3.2. Proof of the theorem
with the subscripts/superscripts $\mathcal{B}_{l}^{\prime},\mathcal{B}_{r}^{\prime}$ to differentiate when the operators correspond to Brownian excursions from $\mathcal{B}_{j}^{\prime}$ to $\mathcal{B}_{k}^{\prime}$, $j,k=l,r$.
We first finish the proof of Theorem 2.12 before going over the proof of Proposition 2.14. When computing the average of Equation (27) with Lemma 2.3, the determinant term is that of [^7] In this proof all occurrences of the Green’s function $G_{D\setminus\eta}$ should be understood as the operator $(G_{D\setminus\eta})_{\delta}^{\delta}$ on $L^{2}(\delta)$, similarly for $G_{D\setminus\eta}$: we drop the subscripts and superscripts to lighten the notation.
(28) $I-G_{D\setminus\eta}(H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})_{\mathcal{B}_{l}^{\prime}}^{\mathcal{B}_{l}^{\prime}}-G_{D\setminus\eta}(H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})_{\mathcal{B}_{r}^{\prime}}^{\mathcal{B}_{r}^{\prime}}-2G_{D\setminus\eta}(H_{\partial D^{\prime}})_{\mathcal{B}_{l}^{\prime}}^{\mathcal{B}_{r}^{\prime}}.$
This can again be understood with another decomposition formula for Brownian loops. One can further decompose the expressions obtained after applying Lemma 1.1 by distinguishing on which side the Brownian path leaves and hits $\delta$:
$G_{D}-G_{D\setminus\eta}=G_{D\setminus\eta}\left((H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})_{\mathcal{B}_{l}^{\prime}}^{\mathcal{B}_{l}^{\prime}}-(H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})_{\mathcal{B}_{r}^{\prime}}^{\mathcal{B}_{r}^{\prime}}-2(H_{\partial D^{\prime}})_{\mathcal{B}_{l}^{\prime}}^{\mathcal{B}_{r}^{\prime}}\right)G_{D}.$
From here, Proposition 1.2 once again provides that the determinant term in the Gaussian integration is equal to
@@
(29) $\exp(\frac{1}{2}((\partial_{\nu_{l}}+\partial_{\nu_{r}})(g_{D^{\prime}}),G_{D}((\partial_{\nu_{l}}+\partial_{\nu_{r}})g_{D^{\prime}}))_{\delta})=\exp(-\frac{1}{2}((\partial_{\nu_{l}}+\partial_{\nu_{r}})g_{D^{\prime}},G_{D}((\partial_{\nu_{l}}+\partial_{\nu_{r}})(g_{D}-g_{D^{\prime}})))_{\delta}),$
given that the gradient of $g_{D}$ is single valued. Now we need to invert $(G_{D})_{\delta}^{\delta}$, i.e. the operator
$|f\mapsto\int_{\delta}G_{D}(\cdot,x)f(x)\absolutevalue{\mathrm{d}x}$
that from a function in $L^{2}(\delta)$ outputs a continuous function in $D$, harmonic in $D\setminus\delta$, with values 0 on $\partial D$ and whose normal derivative jumps by $f(x)$ across $\delta$ at $x$. If one now replaces $g_{D^{\prime}}$ with the multivalued function $g_{D^{\prime},m}$ that is equal to $g_{D}+2n\lambda$ in every branch in $D^{\prime}$, then $g_{D}-g_{D^{\prime},m}$ is single-valued and satisfies the conditions above [^8] In fact the inverse of the Green’s function here is also a Neumann-type operator that computes the difference of the normal derivatives of the harmonic extensions of $f$ on both sides of $\delta$.: it follows that $G_{D}((\partial_{\nu_{l}}+\partial_{\nu_{r}})(g_{D}-g_{D^{\prime},m}))=g_{D}-g_{D^{\prime},m}$. We rewrite Equation (29) as
$\exp(-\frac{1}{2}((g_{D}-g_{D^{\prime},m}),(\partial_{\nu_{l}}+\partial_{\nu_{r}})g_{D^{\prime}})_{\delta})=\exp(\frac{1}{2}\int_{D^{\prime}}\nabla(g_{D^{\prime},m}-g_{D})\cdot\nabla g_{D^{\prime}})$
@@
$\int_{D^{\prime}}\nabla(g_{D^{\prime},m}-g_{D})\cdot\nabla g_{D^{\prime}}=\int_{\partial D^{\prime}}g_{D^{\prime}}\cdot\partial_{\nu}(g_{D^{\prime}}-g_{D})=4\lambda^{2}\int_{\partial D_{l}^{\prime}}\partial_{\nu}(g_{D^{\prime}}-g_{D})=\norm{\nabla g_{D}^{\prime}}_{D^{\prime},\mathrm{reg}}^{2}-\norm{\nabla g_{D}}_{D^{\prime},\mathrm{reg}}^{2}$
given that $g_{D}$ has monodromy $2\lambda$ and $\partial_{\nu_{l}}g_{D}=-\partial_{\nu_{r}}g_{D}$ on $\delta\subset\partial D_{l}^{\prime}$. ∎
We conclude with the proof of Proposition 2.14.
Proof of Proposition 2.14.
The proof starts identically to that of Proposition 2.9, up to the computation of $((u_{\xi}^{\varepsilon})_{\eta},(u_{\xi}^{\varepsilon})_{\eta})_{\nabla(D^{\prime})}$. Write
@@
Similarly to the proof of Proposition 2.9, we assume that $\xi$ is smooth, and compute
@@
$=\int_{\mathcal{B}_{l}^{\prime}}\xi\cdot(H_{\partial D^{\prime}}-H_{\partial D_{l}^{\prime}})_{\mathcal{B}_{l}^{\prime}}^{\mathcal{B}_{l}^{\prime}}\xi+\int_{\mathcal{B}_{r}^{\prime}}\xi\cdot(H_{\partial D^{\prime}}-H_{\partial D_{r}^{\prime}})_{\mathcal{B}_{r}^{\prime}}^{\mathcal{B}_{r}^{\prime}}\xi+2\int_{\mathcal{B}_{l}^{\prime}}\xi\cdot(H_{\partial D^{\prime}})_{\mathcal{B}_{l}^{\prime}}^{\mathcal{B}_{r}^{\prime}}\xi,$
@@
The computation of the term $(\Phi_{\eta},u_{\xi}^{\varepsilon})_{\nabla(D^{\prime})}$ is identical to Equation (22), taking into account that $\mathcal{B}_{l}^{\prime}$ and $\mathcal{B}_{r}^{\prime}$ are seen as distinct parts of the boundary of $D^{\prime}$ (they appear separately when applying the divergence theorem):
$(\Phi_{\eta},u_{\xi}^{\varepsilon})_{\nabla(D^{\prime})}=-\int_{\delta}\xi\cdot\partial_{\nu}h_{\eta}=-\int_{\mathcal{B}_{l}^{\prime}}\xi\cdot\partial_{\nu}h_{\eta}-\int_{\mathcal{B}_{r}^{\prime}}\xi\cdot\partial_{\nu}h_{\eta}=-(\partial_{\nu}g_{D^{\prime}},\xi)_{\mathcal{B}_{l}^{\prime}}-(\partial_{\nu}g_{D^{\prime}},\xi)_{\mathcal{B}_{r}^{\prime}}.\qed$
2.3.3. Corollaries for annular domains
The conditioning on the events $\mathcal{A}_{\mathbf{b}}$ corresponds loosely speaking to two requirements: that the level-line $\eta$ of $\mathbb{P}_{D}^{g_{\mathbf{b}}}$ winds around $\mathcal{B}_{\text{in}}$ (the boundary component containing the target point $y$) a number of times given by $b_{\text{in}}$, and that for each $1\leq j\leq n$ it winds around the other boundary components $\mathcal{B}_{j}$ a number of times given by $b_{j}$. While the latter conditioning is necessary in general, the former turns out to be superfluous: indeed, the monodromy of $g_{\mathbf{b}}$ already fixes the winding of $\eta$ around $\mathcal{B}_{\text{in}}$, in the sense that the level line of $\mathbb{P}_{D}^{g_{\mathbf{b}}}$ almost surely satisfies the property that its winding around $\mathcal{B}_{\text{in}}$ is given by $b_{\text{in}}$. We refer to Appendix A for the proof of this fact. In particular, when $D$ is a conformal annulus (corresponding to $n=0$), there are no other boundary components than those containing the target points $x$ and $y$ and thus no conditioning in the construction of the $\mathrm{SLE}_{4}$ measure. We hence get the following corollary:
Corollary 2.15.
When $D$ is a conformal annulus, we have that
$Z_{D}(x,y)=\sum_{b_{\text{in}}\in\mathbb{Z}}\exp(-\frac{1}{2}\norm{g_{b_{\text{in}}}}_{\nabla(D),\mathrm{reg}}^{2}),$
and $\nu_{D}$ corresponds to the law of a level-line of a compactified GFF $\Phi=\Phi_{D}^{0}+g$, where $g$ is a random multivalued harmonic function in $D$, independent of $\Phi_{D}^{0}$, taking values $g=g_{b_{\text{in}}}$ with probability $Z_{D,b_{\text{in}}}/\sum_{b_{\text{in}}^{\prime}\in\mathbb{Z}}Z_{D,b_{\text{in}}^{\prime}}$.
The law of the winding of the curve under $\nu_{D}$ can then be computed explicitly.
Corollary 2.16.
For $\alpha$ the representative of $(\arg y-\arg x)/(2\pi)$ in $[0,1)$, it holds that
(30) $\nu_{\mathbb{A}_{p}}^{\#}(x,y)[\eta\in\mathcal{A}_{k}]=\frac{Z_{D,k}}{Z_{D}}=\frac{e^{-\frac{\pi^{2}}{2p}\cdot(k+\alpha)^{2}}}{e^{-\frac{\pi^{2}}{2p}\alpha^{2}}\vartheta_{3}(i\frac{\pi}{2p}\alpha|i\frac{\pi}{2p})},\quad k\in\mathbb{Z},$
where $\vartheta_{3}(z|\tau)=\sum_{k\in\mathbb{Z}}e^{\pi ik^{2}\tau}e^{2\pi ikz}$ is a Jacobi theta function.
Proof.
[…]
$\norm{g_{b_{\text{in}}}}_{\nabla(D),\mathrm{reg}}^{2}$ $=\norm{g_{0}}_{\nabla(D),\mathrm{reg}}^{2}+b_{\text{in}}^{2}\norm{f}_{\nabla(D)}^{2}+2b_{\text{in}}(g_{0},f)_{\nabla(D)}$
$=\norm{g_{0}}_{\nabla(D),\mathrm{reg}}^{2}+b_{\text{in}}^{2}\cdot\frac{\pi^{2}}{p}+2b_{\text{in}}(g_{0},f)_{\nabla(D)}$
$\norm{g_{b_{\text{in}}}}_{\nabla(D),\mathrm{reg}}^{2}=\frac{\pi^{2}}{p}\cdot(b_{\text{in}}+\alpha)^{2}+C.\qed$
3. $\mathrm{SLE}_{\kappa}$ as a mixture of boundaries of restriction measures and CLE
@@
We now turn to a more general construction of $\mathrm{SLE}_{\kappa}$ in multiply connected domains for the whole range $\kappa\in(8/3,4]$. Here the constructions and proofs are simpler than in the previous section, but the philosophy of having to separate topological events remains the same. The constructions here are specific to the non-crossing case.
We start by recalling classical facts about restriction measures and CLEs and then state and prove our construction.
3.1. Preliminaries
3.1.1. Restriction measures and Poisson point processes of Brownian excursions
We first recall the definition of one-sided restriction measures, as introduced in [26]. For $D$ a bounded simply connected domain and distinct $x,y\in\partial D$, consider a probability measure $\mathbb{P}$ on closed $R\subset\overline{D}$ such that $R\cap\partial D=(yx)$ and $D\setminus R$ is connected. We say that $\mathbb{P}$ is a one-sided restriction measure with exponent $\alpha$ if there exists $\alpha\geq 0$ such that for all closed $K\subset\overline{D}$ with $K\cap(yx)=\emptyset$ and $D\setminus K$ is simply connected [^9] The $\sigma$-field we consider is generated by the events $R\cap K=\emptyset$, which is the same as the Borel $\sigma$-field generated by the Hausdorff metric on closed subsets $\overline{D}$, see [26, Section 3].,
$\mathbb{P}(R\cap K=\emptyset)=(\Phi_{K}^{\prime}(x)\Phi_{K}^{\prime}(y))^{\alpha}=\left(\frac{H_{\partial(D\setminus K)}(x,y)}{H_{\partial D}(x,y)}\right)^{\alpha},$
where $\Phi_{K}:D\setminus K\to D$ is any conformal equivalence with $\Phi_{K}(x)=x,\Phi_{K}(y)=y$. These restriction measures are conformally invariant and can be constructed in several ways (see for instance [45] for a survey). We will use the construction involving Brownian excursions, which generalises nicely to the multiply-connected setting. Consider $(D,x,y)$ a (non-crossing) admissible domain and a Poisson point process (PPP) of intensity given by $h$ (calling $h$ the intensity constant) times the (infinite) measure of Brownian excursions from $\mathcal{B}_{l}$ in $D$
[…]
where $\mu_{\partial D}(z,w)=H_{\partial D}(z,w)\mu_{\partial D}^{\#}(z,w)$ is the Brownian measure on excursions [^10] We refer to [23, Section 3.3] once again for the precise definition of these measures and the fact that they satisfy conformal invariance and the restriction property, which is all we need for this section. The restriction property states that $\mu_{\partial D}(z,w)$ restricted to curves staying in a subdomain $D^{\prime}\subset D$ agreeing with $D$ near $z$ and $w$ coincides with $\mu_{\partial D^{\prime}}(z,w)$. from $x$ to $y$ in $D$. Denote by $R_{h}^{D,\mathcal{B}_{l}}$ its filling, i.e. the complement of the connected component of $D$ minus a sample of the PPP containing $\mathcal{B}_{r}$. In the case where $D$ is simply connected, $R_{h}^{D,\mathcal{B}_{l}}$ gives rise to the one-sided restriction measure [45]  [^11] The factor of $\pi$ in Equation (31) is due to our convention for the Poisson kernel, in order for $R_{h}^{D,\mathcal{B}_{l}}$ to be a one-sided restriction measure of exponent $h$ as introduced in [26]. Indeed, when $D$ is simply connected, by Equation (8), Equation (32) becomes $\left(\frac{H_{\partial D^{\prime}}(x,y)}{H_{\partial D}(x,y)}\right)^{h},$ i.e. $R_{h}^{D,\mathcal{B}_{l}}$ is a one-sided restriction measure of exponent $h$. The identification of the relation between the intensity constant of the PPP and the value of the parameter of the corresponding restriction measure was also carried out in [28, Theorem 3.7].. The key is the following calculation, which generalises to the multiply-connected setting and that we will make use of in the proof of the Theorem below.
Lemma 3.1.
@@ 3.1.1. Restriction measures and Poisson point processes of Brownian excursions
(32) $\exp(-\frac{h\pi}{4}(\norm{u_{D^{\prime}}}_{\nabla(D^{\prime}),\mathrm{reg}}^{2}-\norm{u_{D}}_{\nabla(D),\mathrm{reg}}^{2})),$
where $u_{D}$ (resp. $u_{D^{\prime}}$) is harmonic with boundary values 1-1 on $\mathcal{B}_{l}$, +1+1 on the rest of $\partial D$ (resp. $\partial D^{\prime}$).
@@
By definition of the PPP and the restriction property satisfied by the excursion measures, the random variable that counts the loops in a realisation of the PPP that exit $D^{\prime}$ is Poisson with parameter
[…]
@@
$\frac{h\pi}{4}(\norm{u_{D^{\prime}}}_{\nabla(D^{\prime}),\mathrm{reg}}^{2}-\norm{u_{D}}_{\nabla(D),\mathrm{reg}}^{2}).\qed$
The same construction can be replicated when replacing $\mathcal{B}_{l}$ with $\mathcal{B}_{\mathbf{b}}^{-}$ for $\mathbf{b}\in\left\{\pm\lambda\right\}^{n}$.
3.1.2. Conformal Loop Ensembles